# Observing radio transients with Phased ALMA: Pulses from the Galactic Centre magnetar

### v2

@@ Acknowledgements.
A significant part of this work was developed during the NRAO REU Chile Internship, conducted at the Joint ALMA Observatory from January to March 2024. We extend our gratitude to AUI/NRAO Chile, D. Rebolledo and P. Cortés for their support. J.V-C gives special thanks to J.C. Fluxá for his support during the NRAO REU Chile Internship application, which ultimately led to this article. We also thank Prof. T. Cassanelli for his valuable feedback. This work was funded by the Max Planck Society through the Max Planck Partner Group led by Prof. M. Cruces.References
- Abbate et al. (2023) Abbate, F., Noutsos, A., Desvignes, G., et al. 2023, Monthly Notices of the Royal Astronomical Society, 524, 2966–2977
- Akiyama et al. (2019) Akiyama, K., Alberdi, A., Alef, W., et al. 2019, The Astrophysical Journal Letters, 875, L2
- Arcus et al. (2024) Arcus, W. R., James, C. W., Ekers, R. D., et al. 2024, arXiv e-prints, arXiv:2408.09351
- Bethapudi et al. (2023) Bethapudi, S., Spitler, L. G., Main, R. A., Li, D. Z., & Wharton, R. S. 2023, Monthly Notices of the Royal Astronomical Society, 524, 3303
@@
- Champion et al. (2020) Champion, D., Cognard, I., Cruces, M., et al. 2020, Monthly Notices of the Royal Astronomical Society, 498, 6044
@@
- Crew et al. (2023) Crew, G. B., Goddi, C., Matthews, L. D., et al. 2023, Publications of the Astronomical Society of the Pacific, 135, 025002
@@
- Gourdji et al. (2019) Gourdji, K., Michilli, D., Spitler, L. G., et al. 2019, The Astrophysical Journal Letters, 877, L19
@@
- Jankowski et al. (2017a) Jankowski, F., van Straten, W., Keane, E. F., et al. 2017a, Monthly Notices of the Royal Astronomical Society, 473, 4436
- Jankowski et al. (2017b) Jankowski, F., van Straten, W., Keane, E. F., et al. 2017b, Monthly Notices of the Royal Astronomical Society, 473, 4436
@@
- Kirsten et al. (2024) Kirsten, F., Ould-Boukattine, O. S., Herrmann, W., et al. 2024, Nature Astronomy, 8, 337
- Kirsten et al. (2021) Kirsten, F., Snelders, M., Jenkins, M., et al. 2021, Nature Astronomy, 5, 414
- Kouveliotou (1999) Kouveliotou, C. 1999, Proceedings of the National Academy of Science, 96, 5351
- Kumar et al. (2023) Kumar, P., Luo, R., Price, D. C., et al. 2023, Monthly Notices of the Royal Astronomical Society, 526, 3652
@@
- Liu et al. (2021) Liu, K., Desvignes, G., Eatough, R. P., et al. 2021, The Astrophysical Journal, 914, 30
- Liu et al. (2019) Liu, K., Young, A., Wharton, R., et al. 2019, The Astrophysical Journal Letters, 885, L10
- Locatelli et al. (2019) Locatelli, N., Ronchi, M., Ghirlanda, G., & Ghisellini, G. 2019, Astronomy & Astrophysics, 625, A109
@@
- Manchester et al. (2005) Manchester, R. N., Hobbs, G. B., Teoh, A., & Hobbs, M. 2005, AJ, 129, 1993
@@
- Mottez et al. (2020) Mottez, F., Zarka, P., & Voisin, G. 2020, Astronomy & Astrophysics, 644, A145
- Nice et al. (2015) Nice, D., Demorest, P., Stairs, I., et al. 2015, Tempo: Pulsar timing data analysis, Astrophysics Source Code Library, record ascl:1509.002
@@
- Pal (2025) Pal, A. 2025, arXiv e-prints, arXiv:2502.11215
- Pearlman et al. (2018) Pearlman, A. B., Majid, W. A., Prince, T. A., Kocz, J., & Horiuchi, S. 2018, The Astrophysical Journal, 866, 160
- Pen & Connor (2015) Pen, U.-L. & Connor, L. 2015, The Astrophysical Journal, 807, 179
@@
- Ransom (2011) Ransom, S. 2011, PRESTO: PulsaR Exploration and Search TOolkit, Astrophysics Source Code Library, record ascl:1107.017
@@
- Serylak et al. (2009) Serylak, M., Stappers, B., Weltevrede, P., et al. 2009, Monthly Notices of the Royal Astronomical Society, 394, 295
- Snelders et al. (2023) Snelders, M. P., Nimmo, K., Hessels, J. W. T., et al. 2023, Nature Astronomy, 7, 1486
- Sobacchi et al. (2024) Sobacchi, E., Iwamoto, M., Sironi, L., & Piran, T. 2024, Escape of fast radio bursts from magnetars
@@
- Torne et al. (2015) Torne, P., Eatough, R. P., Karuppusamy, R., et al. 2015, Monthly Notices of the Royal Astronomical Society: Letters, 451, L50
- Torne et al. (2023) Torne, P., Liu, K., Eatough, R. P., et al. 2023, The Astrophysical Journal, 959, 14
- Wang et al. (2022) Wang, F., Zhang, G., Dai, Z., & Cheng, K. 2022, Nature Communications, 13, 4382
- Witzel et al. (2021) Witzel, G., Martinez, G., Willner, S., et al. 2021, The Astrophysical Journal, 917, 73
@@
- Xu et al. (2023) Xu, J., Feng, Y., Li, D., et al. 2023, Universe, 9, 330
- Yan et al. (2018) Yan, W. M., Wang, N., Manchester, R. N., Wen, Z. G., & Yuan, J. P. 2018, Monthly Notices of the Royal Astronomical Society, 476, 3677
- Yan et al. (2015) Yan, Z., Shen, Z.-Q., Wu, X.-J., et al. 2015, The Astrophysical Journal, 814, 5
- Yao et al. (2017) Yao, J. M., Manchester, R. N., & Wang, N. 2017, The Astrophysical Journal, 835, 29

### v4

@@ Acknowledgements.
A significant part of this work was developed during the NRAO REU Chile Internship, conducted at the Joint ALMA Observatory from January to March 2024. We extend our gratitude to AUI/NRAO Chile, D. Rebolledo and P. Cortés for their support. J.V-C gives special thanks to J.C. Fluxá for his support during the NRAO REU Chile Internship application, which ultimately led to this article. We also thank Prof. T. Cassanelli for his valuable feedback. This work made use of archival data from GMVA 2017 campaign with ALMA (Project Code: 2017.1.00795.V). We extend our gratitude to the P.I. M. Johnson and the team. This work is funded by the Max Planck Society through the Max Planck Partner Group led by Prof. M. Cruces. This work is funded by the Chinese Academy of Sciences President’s International Fellowship Initiative. Grant No. 2026PVC0050).References
- Abbate et al. (2023) Abbate, F., Noutsos, A., Desvignes, G., et al. 2023, MNRAS, 524, 2966–2977
- Akiyama et al. (2019) Akiyama, K., Alberdi, A., Alef, W., et al. 2019, ApJ, 875, L2
- Arcus et al. (2025) Arcus, W. R., James, C., Ekers, R., et al. 2025, PASA, 42, e003
- Bethapudi et al. (2023) Bethapudi, S., Spitler, L. G., Main, R. A., Li, D. Z., & Wharton, R. S. 2023, MNRAS, 524, 3303
@@
- Champion et al. (2020) Champion, D., Cognard, I., Cruces, M., et al. 2020, MNRAS, 498, 6044
@@
- Crew et al. (2023) Crew, G. B., Goddi, C., Matthews, L. D., et al. 2023, PASP, 135, 025002
@@
- Gourdji et al. (2019) Gourdji, K., Michilli, D., Spitler, L. G., et al. 2019, ApJ, 877, L19
@@
- Jankowski et al. (2017a) Jankowski, F., van Straten, W., Keane, E. F., et al. 2017a, MNRAS, 473, 4436
- Jankowski et al. (2017b) Jankowski, F., van Straten, W., Keane, E. F., et al. 2017b, MNRAS, 473, 4436
@@
- Kirsten et al. (2024) Kirsten, F., Ould-Boukattine, O. S., Herrmann, W., et al. 2024, Nat. Astron., 8, 337
- Kirsten et al. (2021) Kirsten, F., Snelders, M., Jenkins, M., et al. 2021, Nat. Astron., 5, 414
- Kouveliotou (1999) Kouveliotou, C. 1999, PNAS, 96, 5351
- Kumar et al. (2023) Kumar, P., Luo, R., Price, D. C., et al. 2023, MNRAS, 526, 3652
@@
- Liu et al. (2021) Liu, K., Desvignes, G., Eatough, R. P., et al. 2021, ApJ, 914, 30
- Liu et al. (2019) Liu, K., Young, A., Wharton, R., et al. 2019, ApJ, 885, L10
- Locatelli et al. (2019) Locatelli, N., Ronchi, M., Ghirlanda, G., & Ghisellini, G. 2019, A&A, 625, A109
@@
- Manchester et al. (2005) Manchester, R. N., Hobbs, G. B., Teoh, A., & Hobbs, M. 2005, ApJ, 129, 1993
@@
- Mottez et al. (2020) Mottez, F., Zarka, P., & Voisin, G. 2020, A&A, 644, A145
- Nice et al. (2015) Nice, D., Demorest, P., Stairs, I., et al. 2015, Tempo: Pulsar timing data analysis, ASCL, record ascl:1509.002
@@
- Pal (2025) Pal, A. 2025, ApJL, accepted, arXiv:2502.11215
- Pearlman et al. (2018) Pearlman, A. B., Majid, W. A., Prince, T. A., Kocz, J., & Horiuchi, S. 2018, ApJ, 866, 160
- Pen & Connor (2015) Pen, U.-L. & Connor, L. 2015, ApJ, 807, 179
@@
- Ransom (2011) Ransom, S. 2011, PRESTO: PulsaR Exploration and Search TOolkit, ASCL, record ascl:1107.017
@@
- Ros et al. (2024) Ros, E., Krichbaum, T. P., Lobanov, A. P., et al. 2024, in Proceedings of the 16th EVN Symposium, ed. E. Ros, P. Benke, S. A. Dzib, I. Rottmann, & J. A. Zensus, 159–168
- Serylak et al. (2009) Serylak, M., Stappers, B., Weltevrede, P., et al. 2009, MNRAS, 394, 295
- Siegel & Fonseca (2025) Siegel, S. & Fonseca, E. 2025, in AAS Meeting Abstracts, Vol. 246, American Astronomical Society Meeting Abstracts, 317.03
- Snelders et al. (2023) Snelders, M. P., Nimmo, K., Hessels, J. W. T., et al. 2023, Nat. Astron., 7, 1486
- Sobacchi et al. (2024) Sobacchi, E., Iwamoto, M., Sironi, L., & Piran, T. 2024, A&A, 690, A332
@@
- Torne et al. (2015) Torne, P., Eatough, R. P., Karuppusamy, R., et al. 2015, MNRAS, 451, L50
- Torne et al. (2023) Torne, P., Liu, K., Eatough, R. P., et al. 2023, ApJ, 959, 14
- Wang et al. (2022) Wang, F., Zhang, G., Dai, Z., & Cheng, K. 2022, Nat. Commun., 13, 4382
- Witzel et al. (2021) Witzel, G., Martinez, G., Willner, S., et al. 2021, ApJ, 917, 73
@@
- Yan et al. (2018) Yan, W. M., Wang, N., Manchester, R. N., Wen, Z. G., & Yuan, J. P. 2018, MNRAS, 476, 3677
- Yan et al. (2015) Yan, Z., Shen, Z.-Q., Wu, X.-J., et al. 2015, ApJ, 814, 5
- Yao et al. (2017) Yao, J. M., Manchester, R. N., & Wang, N. 2017, ApJ, 835, 29